# Weaker Assumptions for Asymmetric Trust

### v1

@@ 7.2 Binding Crusader Agreememt
If all processes with depth $d^{\prime}$ bca-propose a value then every process with depth $d$ eventually bca-decides.
Validity:
If all processes with depth $d^{\prime}$ bca-propose the same input value $v$, then this is the only possible bca-decide output for all processes with depth $d$.
Agreement:
If two processes with depth $d$ bca-decide $v$ and $v^{\prime}$, then $v=v^{\prime}$ or one of the values is $\bot$.
Binding:
Let time $\tau$ be the first time at which there is a process with depth $d$ that bca-decides at time $\tau$. At time $\tau$, there is a value $b\in\{0,1\}$ such that no party with depth $d$ bca-decides 1b1-b in any extension of this execution.
Algorithm 3 presents a solution to this problem for $d^{\prime}=2$ and $d=6$. We prove its correctness in Lemma 7.6. Lemmas 7.37.4, and 7.5 prove helper results which are of use in the proof of the main lemma.
State
$\textit{sentecho2}\leftarrow\textsc{false}$: variable to control the sending of messages
$\textit{approved}\leftarrow\{\}$: array to store potential output values
upon event bca-propose(input) do
send message $\textsc{echo},{\textit{input}}]$ to all
upon receiving a message $\textsc{echo},{\textit{c},\textit{v}}]$ from all processes in some $K\in\mathcal{K}_{i}$ do
send message $\textsc{echo${}^{\prime}$},{\textit{c},\textit{v}}]$ to all
upon receiving a message $\textsc{echo${}^{\prime}$},{\textit{c},\textit{v}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ do
send message $\textsc{echo},{\textit{c}+1,\textit{v}}]$ to all
upon receiving a message $\textsc{echo},{\textit{c},\textit{v}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ do
$\textit{approved}\leftarrow\textit{approved}\cup\{v\}$
if $\neg\textit{sentecho2}$ then
send message $\textsc{echo2},{\textit{v}}]$ to all
$\textit{sentecho2}\leftarrow\textsc{true}$
upon receiving a message $\textsc{echo2},{\textit{v}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ or $|\textit{approved}|>1$ do
if $|\textit{approved}|>1$ then
send message $\textsc{echo3},{\bot}]$ to all
else
send message $\textsc{echo3},{\textit{v}}]$ to all
upon receiving a message $\textsc{echo3},{\textit{v}}]$ from a $Q\in\mathcal{Q}_{i}$ or $(|\textit{approved}|>1\text{ and received {echo3} messages from a }Q\in\mathcal{Q}_{i})$ do
if received $\textsc{echo3},{v}]$ from a $Q\in\mathcal{Q}_{i}$ then
bca-decide($v$)
else
bca-decide($\bot$)
Algorithm 3 Asymmetric binding crusader agreement (code for $p_{i}$)
Lemma 7.3.
Given an execution with faulty processes $F$, a process $p_{i}$ with depth 2, one of its quorums $Q\in\mathcal{Q}_{i}$, and any possible mapping $M:\mathcal{P}\rightarrow\{0,1\}$. There is a value $v\in\{0,1\}$ such that for any other process $p_{j}$ with depth 1, $\exists K\in\mathcal{K}_{j}$ such that $K\subseteq Q\setminus F$ and $\forall p\in K,M(p)=v$
Proof.
We will prove the lemma with the help of a simple algorithm. Each process $p_{i}$ starts the execution with an initial value $v_{i}\in\{0,1\}$, which is sent to all other processes. Afterwards, each process collects values coming from the processes in one of their quorums $Q_{i}\in\mathcal{Q}_{i}$ in a set denoted $V_{i}$.
Suppose that all correct processes execute this algorithm and consider any possible assignment of initial values to the processes, the mapping $M$. We will show that for any process with depth 1 $p_{i}$, there is a value $v$ and a set of correct processes $K\subseteq Q_{i}$ such that all processes in $K$ sent a message with value $v$ and $K$ is a kernel for all processes with depth 1.
$p_{i}$ received messages from $Q_{i}$, let $F$ be the faulty parties in the execution and let $S_{v^{\prime}}\subseteq Q_{i}$ denote the processes in Q which sent value $v^{\prime}$. Suppose $Q_{i}\setminus F$ has no kernel for all other processes with depth 1 such that all of them send the same value $v$. This means that there exists a process with depth 2 $p_{1}$ and a quorum $Q_{1}\in\mathcal{Q}_{1}$ such that $S_{0}\cap Q_{1}=\emptyset$, that is, all correct processes in $Q_{1}$ sent the value 1. This means that the processes in $S_{0}$ cannot form a kernel for all other processes. Similarly, assume that there is a process with depth 1 $p_{0}$ and a quorum $Q_{0}\in\mathcal{Q}_{0}$ such that $S_{1}\cap Q_{0}=\emptyset$. However, if this occurs, $Q_{0}\cap Q_{1}\subseteq F$ which contradicts the consistency property of the quorum system since $p_{0}$ and $p_{1}$ both have depth 1.
This holds for all quorums of all processes with depth 1 and any mapping $M$. Therefore, any quorum contains a kernel with the same value for all other processes with depth 1. This will be of use when proving the termination property of Algorithm 3. ∎
Lemma 7.4.
If all processes with depth 2 start Algorithm 3 with the same input value $v$, all processes with depth 6 will output $v$.
Proof.
The algorithms is used to solve C[9], so we will assume that the execution contains at least one process $p_{8}$ with depth 8. This process has a quorum $Q_{8}$ such that all its members have depth at least 7. Suppose that a process $p_{4}$ with depth 4 receives echo messages with value 1v1-v from one of its quorums $Q_{4}\in\mathcal{Q}_{4}$. Since $Q_{4}\cap Q_{8}\neq\emptyset$, there is at least one process $p_{7}$ with depth 7 that sent an echo message with value 1v1-v towards $p_{4}$. However, $p_{7}$ started with value $v$, so it must have received echo messages with value 1v1-v from one of its quorums $Q_{6}\in\mathcal{Q}_{6}$ (line 3). Since $Q_{6}\cap Q_{7}\neq\emptyset$, there is a process $p_{6}^{\prime}$ with depth 6 that received echo messages with value 1v1-v from one of its kernels $K_{6}^{\prime}\in\mathcal{K}_{6}^{\prime}$ (line 3). Since any kernel for $p_{6}^{\prime}$ must contain at least one process with depth 5, there is a process $p_{5}$ with depth 5 that sent an echo message with value 1v1-v. Using the same logic as before, $p_{5}$ must have received a quorum of echo messages from some $Q_{5}\in\mathcal{Q}_{5}$. Since $Q_{5}\cap Q_{7}\neq\emptyset$, some process $p_{6}^{\prime\prime}$ with depth 6 must have received echo messages with value 1v1-v from one of its kernels. Following this reasoning, we reach an infinite sequence of processes with depth 6 that must have received echo messages with value 1v1-v from one of their kernels. But since no process with depth 2 has initial value 1v1-v, we can conclude that this is not possible and therefore no process with depth 4 receives a quorum of echo messages with value 1v1-v.
We have established that all processes with depth 4 will only receive a quorum of echo messages with value $v$. Therefore, they will send a $\textsc{echo2},{v}]$ message (line 3 and all processes with depth 5 will receive a quorum of this messages, causing them to send a $\textsc{echo3},{v}]$ message. Finally, all processes with depth 6 will receive a quorum of these messages and deliver value $v$ (line 3).
Lemma 7.5.
If a process with depth 5 sends a $\textsc{echo3},{v}]$ message in Algorithm 3, no other process with depth 5 will send a $\textsc{echo3},{1-v}]$ message.
Proof.
Assume that process $p_{5}$ with depth 5 is the first process with depth 5 to send a echo3 message and does so with value $v$. To do this, it must have received echo2 messages with value $v$ from one of its quorums $Q_{5}\in\mathcal{Q}_{5}$(line 3). Imagine that there is another process $p_{5}^{\prime}$ with depth 5 that sends a $\textsc{echo3},{1-v}]$ message. This requires receiving $\textsc{echo2},{1-v}]$ messages from one of its quorums $Q_{5}^{\prime}\in\mathcal{Q}_{5}^{\prime}$. Since $(Q_{5}\cap Q_{5}^{\prime})\setminus F\neq\emptyset$, and a nonfaulty party only sends echo2 messages with one value, $p_{5}^{\prime}$ could not have sent a $\textsc{echo2},{1-v}]$ message. ∎
Lemma 7.6.
Algorithm 3 solves asymmetric binding crusader agreement with $d=6$ and $d^{\prime}=2$.
Proof.
Termination:
We assume that all processes with depth $d^{\prime}=2$ start the execution. Upon starting they send a a $\textsc{echo},{1,v}]$ message with their respective initial value $v$ (line 3). Since all processes with depth $d^{\prime}+1$ have a quorum formed by processes of depth $d^{\prime}$, they will eventually receive $\textsc{echo},{1,*}]$ messages from one of their quorums. Some of these messages will have $v=0$ and others $v=1$. By Lemma 7.3 we know that there is a value $v\in\{0,1\}$ such that all processes in the quorum which sent $v$ are also a kernel for all other processes with depth 2. Therefore, all processes with depth 2 will receive echo messages with the same value from one of their kernels.
This causes them to send a $\textsc{echo${}^{\prime}$},{1,v}]$ message (line 3). Since all processes with depth 2 will send this message, all processes with depth 3 are guaranteed to receive $\textsc{echo${}^{\prime}$},{1,v}]$ from one of their quorums. This triggers them to send a $\textsc{echo},{2,v}]$ message (line 3). Since all depth $d^{\prime}+1=3$ processes send this message, all depth $d^{\prime}+2$ messages will receive a quorum of $\textsc{echo},{2,v}]$ messages, triggering the sending of a $\textsc{echo2},{v}]$ message if none has been sent so far (line 3). Consequently, all processes with depth $d^{\prime}+3$ will receive a quorum of $\textsc{echo2},{v}]$ messages, send a $\textsc{echo3},{v}]$ message, and all processes with depth $d^{\prime}+4$ will receive a quorum of $\textsc{echo3},{v}]$ messages. This will trigger them to deliver a value, $v$ or $\bot$, depending on the messages received. This shows that all processes with depth 6 will deliver a value.
Agreement: Suppose processes $p_{0}$ and $p_{1}$, both with depth $d$, output values 0 and 1 respectively. This means that $p_{0}$ received $\textsc{echo3},{0}]$ messages from one of its quorums $Q_{0}\in\mathcal{Q}_{0}$ and $p_{1}$ received $\textsc{echo3},{1}]$ messages from one of its quorums $Q_{1}\in\mathcal{Q}_{1}$. If $d\geq 1$, quorum consistency guarantees that $Q_{0}\cap Q_{1}\nsubseteq F$. Let $p_{k}$ be one correct process in $Q_{0}\cap Q_{1}$. Since $p_{0}$ and $p_{1}$ delivered different values, $p_{k}$ must have sent $\textsc{echo3},{0}]$ to $p_{0}$ and $\textsc{echo3},{1}]$ to $p_{1}$. But since $p_{k}$ is correct and the algorithm states that each process sends only one echo3 message this is not possible.
Validity: Follows from Lemma 7.4.
Binding: Consider the first process $p_{6}$ with depth $d=6$ that sees Echo3 messages from one of its quorums $Q_{6}\in\mathcal{Q}_{6}$. There are two cases, if a process with depth 5 sent a non-$\bot$ message then we have binding from Lemma 7.5. Otherwise, all processes in $Q_{6}\setminus F$ sent a $\textsc{echo3},{\bot}]$ message. From Lemma 7.1 we know that $Q_{6}\setminus F$ has a kernel for any other process $p_{6}^{\prime}$ with depth 6. Therefore, it is not possible for $p_{6}^{\prime}$ to see a quorum of Echo3 messages for any non-$\bot$ value.
We have shown that, if we take $d^{\prime}=2$ and $d=6$, the algorithm satisfies all required properties. Therefore, Algorithm 3 solves asymmetric binding crusader agreement with $d=6$ and $d^{\prime}=2$. ∎
7.3 Consensus
Definition 7.3 introduces the concept of depth to the asymmetric consensus problem. We propose a consensus algorithm using the round-based paradigm of Abraham et al. [1]. Each round consists of a call to the binding crusader agreement and common coin primitives presented earlier. If the values obtained from the two primitives match, then the processes deliver it; otherwise, another round is executed.
We make a small addition to the common coin and binding crusader agreement calls. When one of these sub-protocols is triggered, it is done so with a round tag. This is used to keep track of the round that is running and does not interfere with any details of the sub-protocols.
Definition 7.3 (Depth-characterized asymmetric binary Byzantine consensus).
A protocol for Byzantine consensus, shortened as C[d], defined through the events c-propose(val) and c-decide(val), satisfies:
Probabilistic termination:
@@ 7.3 Consensus
If all correct processes c-propose the same value val, then all processes with depth $d$ must c-decide val.
@@
A difference with respect to symmetric round-based protocols is that not all processes will finish all the sub-protocol calls. Only processes with depth 6 can complete the crusader agreement, and only those with depth 7 also finish the subsequent common coin. If we take a naive approach, executing rounds one after another, where only those processes that completed the previous round participate in the next, the required depth increases with the number of rounds. Depth 7 is required to finish the first round, 14 for the second round, 21 for the third and so on.
@@
To fix this, we introduce a revival mechanism, which lets any process with depth at least 2 rejoin future rounds if it ever gets stuck. Without this element, after each round, liveness could only be guaranteed for processes with higher and higher depths. Since such a protocol can have executions with infinite length, we would only be able to provide guarantees for processes with infinite depth, which goes against our objectives here.
The revival mechanism is similar to the quorum amplification technique proposed by Bracha [7]. Each process that finishes the pipeline broadcasts a revive1 message. If termination for a round is guaranteed for processes with depth $7$ but there is also a process with depth $9$, Lemma 7.3 implies that there will be a kernel for all processes with depth 1 formed by processes that finish the pipeline with the same value. Upon receiving $\textsc{revive1},{v,round}]$ messages from a kernel, each process sends a $\textsc{revive2},{v,round}]$ message. We know that all processes with depth at least 1 will send this message. Afterwards, upon receiving revive2 messages from all processes in a quorum, the process will go back to the beginning of the round pipeline. This will happen for all processes with depth 2. Therefore, if all processes with depth $7$ finish the pipeline and we also have at least one process with depth $9$, we can revive all processes with depth 2. This is the reason why our protocol solves C[9] but not C[7], at least one process with depth 9 is needed to make the revival mechanism work.
@@
A process may execute multiple rounds in parallel. If the revival mechanism signals a process to start the next round, while the current round is still ongoing, the process proceeds with both executions simultaneously. This is enabled by the usage of arrays (v, continue) to store the process states separately for each round. This design ensures that the process can continue sending messages for every round it has initiated, which may be needed by other processes in the system.
@@
Algorithm 4 presents a solution to the asymmetric consensus problem for all processes with depth 9. It relies on the binding crusader agreement and common coins previously introduced. Lemma 7.13 proves that it satisfies the properties described in Definition 7.3. Since we want to prove the properties for all processes with depth 9, in the proofs, we will assume that there is at least one such process.
Lemma 7.7.
If a process with depth at least 7 has the variable continue[round+1] set to true (i.e. start the round execution) in Algorithm 4, then at least one processes with depth at least 7 has continue[round] set to true (i.e., finishes the previous round execution).
@@
We are solving C[9], therefore, we will assume the existence of at least one process $p_{9}$ with depth 9.
Assume that process $p_{7}$ sets continue[round+1] to true. There are two ways in which this can happen. It can either receive a quorum of $\textsc{revive2},{val,round+1}]$ messages (line 4) or obtain an output from the common coin (line 4. In the second case, it is clear that at least one process with depth 7 has $\textit{continue}[\textit{round}]=\textsc{true}$, as $p_{7}$ itself satisfies this. This follows from the fact that for $p_{7}$ to set the variable to true, it must have obtained an output from the common coin (line 4) and before that from the binding crusader agreement (line 4). To obtain a bca-decide, it must have first called bca-propose, which only happens if v[continue] is true.
In the other case, if it received $\textsc{revive2},{val,round+1}]$ messages from a quorum $Q_{7}\in\mathcal{Q}_{7}$, we know that all members of $Q_{7}$ received a kernel of $\textsc{revive1},{val,round+1}]$ messages. Let $Q_{9}\in\mathcal{Q}_{9}$ be the quorum of $p_{9}$ such that all its members have depth at least 8 and denote by $p_{8}$ a process that belongs to $Q_{7}\cap Q_{9}$. Since $p_{8}$ has depth 8 and it received revive1 messages from one of its kernels $K_{8}\in\mathcal{K}_{8}$, we know that at least one process $p_{7}^{\prime}$ with depth 7 sent a $\textsc{revive1},{val,round+1}]$ message. By the argument used in the previous paragraph, we know that a process only sends this message if it has $\textit{v}[\textit{round}]=\textsc{true}$. This proves the lemma. ∎
Lemma 7.8.
If all processes with depth at least 2 have the variable continue[round] set to true (i.e. start the round execution) in Algorithm 4, then all processes with depth at least 7 will set continue[round+1] to true (i.e., finish the round execution).
@@
A process with depth at least 7 sets [round+1] to true upon finishing the execution of a round (line 4). From the termination properties of the binding crusader agreement and the common coin we know that if all processes with depth 2 start a round, then all processes with depth 6 will finish the binding crusader agreement (i.e., reach line 4) and all processes with depth 7 will finish the common coin (i.e., finish line 4). All correct processes that finish the common coin call will also finish the round, as there is no other point in which they can be stopped between lines 4 and 4. Therefore, all processes with depth 7 will set continue[round+1] to true and continue executing the next round. This also implies that all processes with depth at least 7 will send a $\textsc{revive1},{v,round}]$ message for some value v. ∎
Lemma 7.9.
If all processes with depth at least 7 set continue[round+1] to true in line 4 of Algorithm 4, then all processes with depth 2 will eventually have $\textit{continue}[\textit{round}+1]$ also set to true.
@@
We are solving C[9], therefore, we will assume the existence of at least one process $p_{9}$ with depth 9. Since all processes with depth at least 7 send a revive1 message for round $\textit{round}+1$, $p_{9}$ will receive such messages from all processes in one of its quorums $Q_{9}\in\mathcal{Q}_{9}$. By Lemma 7.3, we know that there will be a value val and a subset of correct processes $K\subseteq Q_{9}$ such that all processes in $K$ sent $\textsc{revive1},{val,round+1}]$ and $K$ is a kernel for all processes with depth 1. Therefore, all processes with depth 1 will send $\textsc{revive2},{val,round+1}]$ message (line 4) and all processes with depth 2 will receive a quorum of such messages. This triggers them to enter round $r+1$ by setting the variable $\textit{continue}[\textit{round}+1]$ to true and the variable $\textit{v}[\textit{round}+1]$ to the value from the kernel. ∎
Lemma 7.10.
If all processes with depth at least 7 send a $\textsc{revive1},{val,round}]$ message with the same value val, then all processes with depth at least 2 will receive a $\textsc{revive2},{val,round}]$ message from one of their quorums and will not receive $\textsc{revive2},{1-val,round}]$ messages from one of their quorums.
Proof.
Since we are providing a solution for C[9] we assume that there exists at least one process $p_{9}$ with depth 9. Since all processes with depth 7 send this message, $p_{9}$ will receive $\textsc{revive1},{val,round}]$ from at least one of its quorums $Q_{9}\in\mathcal{Q}_{9}$. From Lemma 7.1, we know that $Q_{9}\setminus F$ is a kernel for all processes with depth at least 1, causing all such processes to send a $\textsc{revive2},{val,round}]$ message (line 4). All processes with depth at least 2 will receive this message from one of their quorums.
We now show that no process with depth at least 2 will receive $\textsc{revive2},{1-val,round}]$ messages from all processes in one of its quorums. Suppose that one such process $p_{2}$ receives these messages from a quorum $Q_{2}\in\mathcal{Q}_{2}$. Since $Q_{2}\cap Q_{9}\neq\emptyset$, there is at least one process $p_{8}$ with depth 8 that sent a $\textsc{revive2},{1-val,round}]$ message. $p_{8}$ would have done this after receiving $\textsc{revive1},{1-val,round}]$ from one of its kernels $K_{8}\in\mathcal{K}_{8}$. From the definition of depth we know that $p_{8}$ has a quorum $Q_{8}\in\mathcal{Q}_{8}$ such that all its members have depth at least 7. Since $K_{8}\cap Q_{8}\neq\emptyset$, this means that a process with depth 7 must have sent a $\textsc{revive1},{1-val,round}]$ message, but by the lemma assumptions this is not possible.
Since all processes with depth at least 2 receive a quorum of $\textsc{revive2},{val,round}]$ messages, but no quorum of $\textsc{revive2},{1-val,round}]$ messages, we have proven the lemma. This also implies that they will set v[round] to the value val and continueround] to true (line 4), enabling them to continue execution in round round. ∎
Lemma 7.11.
If a process with depth at least 2 receives $\textsc{revive2},{v,round}]$ messages from one of its quorums, then at least one process with depth at least 7 sent a $\textsc{revive1},{v,round}]$ message.
Proof.
As we did before, we assume that there exists at least one process $p_{9}$ with depth 9. Suppose a process $p_{2}$ with depth 2 received $\textsc{revive2},{v,round}]$ messages from one of its quorums $Q_{2}\in\mathcal{Q}_{2}$. Since $Q_{2}\cap Q_{9}\neq\emptyset$, this means that at least one process $p_{8}$ with depth 8 sent a $\textsc{revive2},{v,round}]$ message. Following the algorithm, this only happens if $p_{8}$ received $\textsc{revive1},{v,round}]$ messages from one of its kernels $K_{8}\in\mathcal{K}_{8}$ (line 4). Since $p_{8}$ has at least one quorum composed only of processes with depth 7, $K_{8}$ contains one such process. Therefore, a process with depth 7 sent a $\textsc{revive1},{v,round}]$ message. ∎
Lemma 7.12.
If all processes with depth at least 2 have $\textit{v}[\textit{round}]$ set to the same value val in Algorithm 4, then all processes with depth 9 will c-decide val within a constant expected number of rounds.
Proof.
From Lemma 7.4 we know that all processes with depth 6 will obtain the same value val as a result of the binding crusader agreement. We also know that all processes with depth at least 7 will output a common coin value c. If the random value of the coin c matches val, then all processes with depth 9 will decide val in this round. Otherwise, by Lemma 7.10, all processes with depth 2 will start the next round with value val. This continues until a round is reached in which $\textit{c}=\textit{val}$, which triggers all processes with depth at least 9 to c-decide val. This happens with probability frac12frac{1}{2} in each round. This follows a geometric distribution and will therefore occur within an expected constant number of rounds. ∎
Lemma 7.13.
@@
Probabilistic Termination: If in a round $r$ all processes with depth 6 output the same value, $\bot$ or val, in their binding crusader agreement, by Lemma 7.12, the execution will finish in an expected constant number of rounds.
Consider the first round of the execution. if all processes with depth at least 6 output the same value, $\bot$ or val, in their binding crusader agreement, by Lemma 7.12 the execution will finish in an expected constant number of rounds.
Otherwise, by the binding property of the crusader agreement, the adversary is bound to some non-$\bot$ value $v$ by the time the first process with depth 6 finishes the crusader agreement. By the unpredictability property of the common coin, this must happen before the coin value c is revealed. Therefore, with probability $\frac{1}{2}$, $\textit{val}=\textit{c}$ and all processes with depth at least 7 set v[round+1] to value val. Then, by applying Lemma 7.10 and then Lemma 7.12, we conclude that within a constant expected number of rounds, a round will be reached where all processes with depth 9 will decide. Otherwise, some processes with depth 7 will set v[round+1] to value val and others will set it to the value c of the coin. By Lemma 7.8 we know that all of them will set continue[round+1] to true. Applying Lemma 7.9 we know that eventually, all processes with depth at least 2 will set continue[round+1] to true and continue the execution in the next round.
The next round presents the same situation, where some processes with depth at least 2 start round+1 with value 0 and others do it with value 1. This will occur until the value of the coin matches the output of the binding crusader agreement, which will also take a constant expected number of rounds.
Agreement: Consider the first round $r$ in which a process decides a value $b$. The value $b$ must be the output from both the binding crusader agreement and the common coin (line 4). Since the value decided must match the coin, any process that decides in round $r$ will also decide $b$. All processes that complete this round will do so with the value $b$ stored in the variable v. Processes that obtain $\bot$ from the crusader agreement will set v to be the value of the coin (line 4), while all other processes will set v to be the output of the crusader agreement (line 4). Therefore, all processes with depth 7 finish the round with $b$ stored in the variable $v$, and when processes with depth 2 are revived, following Lemma 7.10, they do so with value $b$ in their variable v.
We know from Lemma 7.11 that a process with depth 2 will not start round round+1 until at least one process with depth 7 has finished the previous round. Therefore, all processes with depth 2 start the next round and do so with value $b$. Lemma 7.4 shows that all processes with depth 6 will output $b$ from the crusader agreement, and as such no process will decide 1b1-b in the round. If a decision takes place in this round it will be with value $b$, otherwise, all processes with depth 2 start the next round with value $b$. This continues for all subsequent rounds and as a consequence agreement holds.
Validity: We assume that all correct processes (i.e., those with depth 0) start the execution with the same value. Therefore they will set $\textit{v}[1]$ to val. By Lemma 7.12 we know that all processes with depth 9 will c-decide val.
State
$\textit{v}[\ ]$: array that stores input values for each round, initially empty
$\textit{continue}[\ ]\leftarrow[\textsc{false},\dots,\textsc{false}]$: array for flow control, initially all false
upon event c-propose(value) do
v[1] $\leftarrow$ value
$\textit{continue}[1]\leftarrow\textsc{true}$
upon exists round such that $\textit{continue}[\textit{round}]$ do
continue[round] $\leftarrow\textsc{false}$
bca-propose(v[round]) with tag round// calls the BCA protocol
upon bca-decide(value) with tag round do // after finishing the BCA protocol
v[round+1] $\leftarrow$ value
release-coin() with tag round // call the common coin protocol
@@
if c = v[round+1] then
c-decide(v[round+1])
else
v[round+1] $\leftarrow$ c
send message $\textsc{revive1},{\textit{v}[\textit{round}+1],\textit{round}}]$ to all $p\in\mathcal{P}$
$\textit{continue}[\textit{round}+1]\leftarrow\textsc{true}$
upon receiving $\textsc{revive1},{\textit{v${}^{\prime}$},\textit{round}}]$ from all processes in some $K\in\mathcal{K}_{i}$ do
send message $\textsc{revive2},{\textit{v${}^{\prime}$},\textit{round}}]$ to all $p\in\mathcal{P}$
upon receiving $\textsc{revive2},{\textit{v${}^{\prime}$},\textit{round}}]$ from all processes in some $Q\in\mathcal{Q}_{i}$ do
$\textit{v}[\textit{round}]\leftarrow\textit{v}^{\prime}$
$\textit{continue}[\textit{round}]\leftarrow\textsc{true}$
Algorithm 4 Asymmetric binary consensus (code for $p_{i}$)
8 Conclusions and Future Work
Prior to this work, solutions to problems such as reliable broadcast and consensus typically relied on the existence of a guild. Theorem 4.1 shows that this assumption effectively reduces to the existence of a symmetric system, constructed from the original assumptions, that provides equivalent or stronger guarantees. In other words, existing solutions are not well-suited for genuinely asymmetric settings.
To overcome this limitation, we introduced a more fine-grained way to characterize asymmetry through the notion of depth. We proved that reliable broadcast can be solved for all processes with depth at least 3, and consensus for processes with depth at least 9. This significantly relaxes the requirements compared to prior work by Alpos et al. [2], which established feasibility only for processes with infinite depth.
The exact minimum depth required for solving asymmetric reliable broadcast and consensus remains an open problem. For reliable broadcast, the lower bound is either 2 or 3; for consensus, it lies between 1 and 9. Determining these thresholds is crucial, as they capture the precise necessary and sufficient conditions for solving problems under asymmetric trust. Previous work [2, 17] established sufficiency but left necessity unproven.
Finally, an important direction for future research is to explore whether the concepts and techniques introduced here extend to other asymmetric trust models [17, 19].

### v2

@@ 7.3 Consensus
If all correct processes c-propose the same value val, then no processes with depth $d$ c-decides $1-\textit{val}$.
@@
A difference with respect to symmetric round-based protocols is that not all processes will finish all the sub-protocol calls. Only processes with depth 8 can complete the graded agreement, and only those with depth 9 also finish the subsequent common coin. If we follow a naive approach, in which only those processes that completed the previous round participate in the next, the required depth increases with the number of rounds.
@@
We address this fact with the strong termination property of binding grading agreement, which guarantees liveness, even when not every process starts the execution of the protocol. Without this element, after each round, liveness could only be guaranteed for processes with higher and higher depths. Since such a protocol can have executions with infinite length [DBLP:journals/jacm/FischerLP85], we would only be able to provide guarantees for processes with infinite depth, which goes against our objectives here.
@@
A process may execute multiple rounds in parallel. This is enabled by the usage of arrays (send, sent) to store the process states separately for each round. This design ensures that the process can continue sending messages for every round it has initiated, which may be needed by other processes in the system.
@@
Algorithm 4 presents a solution to the asymmetric consensus problem for all processes with depth 9. It relies on the binding graded agreement and common coins previously introduced. Theorem 5 proves that it satisfies the properties described in Definition 12. Lemmas 1513, and 14 show helper results that are used in the proof of the theorem. Since we want to prove the properties for all processes with depth 9, in the proofs, we will assume that there is at least one such process.
State
$\textit{bga-out}\leftarrow[\ ]$: array that stores bga decisions for each round, initially empty
${send}\leftarrow[\textsc{false},\textsc{false}]^{\infty}$: $send[r][v]$ indicates if $v$ should be proposed for round $r$
${sent}\leftarrow[\textsc{false},\textsc{false}]^{\infty}$: $sent[r][v]$ indicates if $v$ has been proposed for round $r$
upon event c-propose(value) do
${send[{value}]}[1]\leftarrow\textsc{true}$
upon exists round and value where $send{[\textit{round}][value]}$ and $\neg sent[[\textit{round}]{value}]$ do
$sent[round][{value}]\leftarrow\textsc{true}$
bga-propose(value) with tag round
upon bga-decide((value, grade)) with tag round do // after finishing the BGA protocol
bga-out[round] $\leftarrow$ (value, grade)
release-coin() with tag round // call the common coin protocol
$(\textit{value},\textit{grade})\leftarrow\textit{bga-out}[round]$
if $grade<2$ then
$send[round+1][{c}]\leftarrow\textsc{true}$
if $grade>0$ then // $value\neq\bot$
$send[round+1][{value}]\leftarrow\textsc{true}$
if $grade=2$ and $\textit{c}=value$ then
c-decide(value)
Algorithm 4 Asymmetric binary consensus (code for $p_{i}$)
Lemma 13.
If for all processes with depth at least 3 a kernel of correct processes bga-propose the same value $v$ with tag round in Algorithm 4, then for some value $v^{\prime}$: all processes with depth at least $9$ bga-propose $v^{\prime}$ with tag round+1.
@@
Assume that for all processes with depth at least 3 a kernel of correct processes bga-propose the same value $v$ with tag round (line 4), then by the strong termination property of BGA[2,8], all processes with depth $8$ bga-decide on a value and a grade and thus invoke CC[8,9] with tag round (line 4). Then all processes with depth $9$ receive output-coin for some value $c\in\{0,1\}$ (line 4) due to termination of CC[8,9]. These processes evaluate both the if statement in line 4 and line 4. Due to graded agreement of BGA[2,8] and the overlap on $g=1$ for the two conditions: if a process with depth $9$ has grade $g=2$, then all processes with depth 9 evaluate line 4 to true (and propose the same value as $g>0$). Likewise, if a process with depth $9$ has grade $g=0$, then all processes with depth $9$ evaluate line 4 to true (and propose the same value due to the matching property of $\textsl{CC}[8,9]$). ∎
Lemma 14.
If in Algorithm 4, all processes with depth at least $9$ bga-propose $v$ with tag round for the same value $v$, then these processes form a kernel for all processes with depth 3.
@@
Assume the existence of at least one process $p_{10}$ with depth 10. Assume that all processes with depth at least 9 bga-propose $v$ with tag round for the same value $v$. Consider a quorum for $p_{10}$ consisting of a subset of these depth $9$ processes. By Lemma 8 these are a kernel for all processes with depth $1$.
Lemma 15.
The matching property of CC[8,9] holds for all processes with depth $2$.
@@
This follows directly from the proof of Theorem 3, as the proof of the matching property does not rely on the level of a correct process. ∎
Theorem 5.
@@
Validity: We assume that all correct processes (i.e., those with depth 0) start the execution with the same value val. Therefore, they $\textsl{bga-propose}(\textit{val})$ and have no path to $\textsl{bga-propose}(\textit{1-val})$. The validity property of binding graded agreement ensures that every process of depth 88 $\textsl{bga-decides}(\textit{val,2})$. Every process with depth $9$ receives the output of the coin and sets $\textit{propose}[\textit{round}][\textit{val}]$ to true, regardless of the value of the coin. Thus, no process with depth 2 $\textsl{bga-proposes}(1-\textit{val})$ and the validity of binding graded agreement, iterating over the same argument, ensures that no process with depth 11 $\textsl{c-decide}(1-\textit{val})$.
@@
8 Conclusion and Future Work
Prior to this work, solutions to problems such as reliable broadcast and consensus relied on guild assumptions. Theorem 1 shows that this assumption effectively reduces to the existence of a symmetric system, constructed from the original assumptions, that provides equivalent or stronger guarantees. In other words, existing solutions are not well-suited for genuinely asymmetric settings.
To overcome this limitation, we introduced a more fine-grained way to characterize asymmetry through the notion of depth. We proved that reliable broadcast can be solved for all processes with depth at least 3, and consensus for processes with depth at least 9. This significantly relaxes the requirements compared to prior work by Alpos et al. [DBLP:journals/dc/AlposCTZ24], which established feasibility only for processes with infinite depth (i.e., guild members).
The exact minimum depth required for solving asymmetric reliable broadcast and consensus remains an open problem. For reliable broadcast, the lower bound is either 2 or 3; for consensus, it lies between 1 and 9. Determining these thresholds is crucial, as they capture the precise necessary and sufficient conditions for solving problems under asymmetric trust. Previous work [DBLP:journals/dc/AlposCTZ24, DBLP:conf/wdag/LiCL23] established sufficiency but left necessity unproven.
Finally, an important direction for future research is to explore whether the concepts and techniques introduced here extend to other asymmetric trust models [DBLP:conf/wdag/LiCL23, DBLP:conf/wdag/LosaGM19].